# Frequency effects in Linear Discriminative Learning

### v1

@@ 6.2 High-dimensional modeling
A model using FIL obtained an average accuracy@1 of 22.3%, which is clearly superior to the 9.9% obtained for the lower-dimensional model. The upper panel of Figure 6 presents the predicted accuracy curves for the high-dimensional FIL model in red, and the low-dimensional FIL model in black. We see a rise in accuracy for lower frequencies.
@@
Figure 6: Predicted accuracy@1 as a function of log frequency for a high-dimensional representations of form and meaning (red line). The black line shows the predicted accuracy based on the low-dimensional model, for comparison. The light blue dots represent the target correlations in the high-dimensional model. The small dataset refers to the dataset with 2,638 word forms based on Ernestus and Baayen, (2003), the large dataset to the dataset created from the DLP (Brysbaert and New,, 2009) including 13,669 word forms.
6.3 Increasing the dataset size
@@
Having established the relationship between EL, WHL and FIL, we can also investigate how EL and FIL fare when the modelled dataset is significantly larger. To this end we used all 13,669 words from the Dutch Lexicon Project (DLP; Keuleers et al.,, 2010) for which we were able to obtain fasttext embeddings, and modelled the dataset using trigram representations (resulting in a form vector dimensionality of 4,678) as well as frequencies extracted from CELEX (as included in the DLP). We found that accuracies in general were clearly lower: For EL, the accuracy@1 was 67.8% and for FIL it was 5.1%. Qualitatively, the differences between EL and FIL are therefore similar (see also Figure 6), but our simple linear mappings clearly perform less well with very large datasets, especially when taking into account frequency.
6.4 Frequency-weighted accuracy
@@
A final question is whether the way we have been evaluating accuracy so far is reasonable. In the average accuracy@1 each word's accuracy contributes the same, that is, we have effectively calculated accuracy across all word types in our corpus. However, it can be argued that comprehending high frequency words is much more important than low frequency words, and that we should be calculating accuracy across word tokens instead. Practically this would mean that if we would go through a corpus of written English, instead of counting how many unique words our model is able to comprehend, we would count how many of all of the encountered word tokens are understood correctly.
@@
Following this line of argumentation we also provide frequency-weighted accuracies in Table 1. For instance, the word with the highest frequency in the Dutch Lexicon Project (large dataset) is de (eng. the). Since this word accounts for 7% of all word tokens in the Dutch Lexicon Project (calculated by summing up the frequencies of all word types), it also contributes 7% to our frequency-weighted accuracy measure. There are also 349 words with a frequency of 0 in CELEX (2.6% of all word types) which accordingly do not contribute to the frequency-weighted accuracy at all. We find that with this method, the results flip in comparison with average accuracy@1: Generally, FIL and WHL based on untransformed frequencies perform the best, followed by methods based on log-transformed frequencies, while EL clearly performs the worst. For example, for the high-dimensional simulations on the large dataset, EL has a frequency-weighted accuracy@1 of 43.7%, while FIL achieves 79.8%.
@@ 7.1 Visual word recognition in Dutch
In order to assess the possibilities offered by FIL-based modeling (using untransformed frequencies and the high-dimensional form and meaning vectors) for predicting behavioral measures of lexical processing, we return to the large dataset of reaction times to Dutch words described in Section 6.3 (13,669 words represented as trigrams with a form dimensionality of 4,678, and semantics represented using 300-dimensional fasttext vectors).
Figure 7 presents the partial effects according to three Gaussian Location-Scale Generalised Additive Models (Wood,, 2011, GAMs,) fitted to the response latencies in the Dutch Lexicon Project. Response latencies were inverse transformed (-1000/RT) in order to avoid marked non-normality in the residuals. (Effects for the untransformed RTs are very similar in shape, but confidence intervals are not reliable due to the marked departure from normality of the residuals). The left-hand panels present the partial effects for the mean (in blue), the right panels the partial effects for the variance (in green, on the $\log(\sigma-0.01)$ scale). The upper panels pertain to a GAM predicting RT from log frequency (AIC -16384.8). Mean and variance decrease non-linearly with increasing frequency. The smooth for the mean shows the kind of non-linearity that is typically observed for reaction time data (see, e.g., Baayen,, 2005; Miwa and Baayen,, 2021): the effect of log frequency levels off strongly for high-frequency words and to a lesser extent also for low-frequency words. The partial effect of the variance is less wiggly and decreases as frequency increases. This decrease in variability for increasing frequency has at least two possible sources. First, high-frequency words are known to all speakers, whereas low-frequency words tend to be specialized and known to smaller subsets of speakers. Second, more practice, as in the case of high-frequency words, typically affords reduced variability (see, e.g., Tomaschek et al.,, 2018).
@@
Figure 7: Partial effects for mean (left, confidence intervals in blue) and variance (right, confidence intervals in green, the y-axis is on the $\log(\sigma-0.01$ scale), for Gaussian Location-Scale GAMs predicting reaction times from log frequency (upper panels), from $1-r$ based on FIL (center panels), and from $1-r$ based on EL (bottom panels). The vertical red lines represent the 0%, 25%, 50%, 75%, and 100% percentiles.
@@
A model-based measure that we expected to correlate with reaction time is the proximity of a word's predicted semantic vector to its corresponding gold standard vector. The more similar the two vectors are, the better a word's meaning is reconstructed from its form. In other words, the more effective a word form is in approximating its meaning, the more word-like it is and the faster a lexicality decision can be executed. We used the correlation $r$ of $\hat{\mathbf{s}}$ and $\mathbf{s}$ as a measure of semantic proximity. Since for large $r$, lexical decision times are expected to be short, whereas for small $r$ long decision times are more likely, we took $1-r$ as a measure that we expect to enter into a positive correlation with RT. This measure, when based on FIL, has a density that is roughly symmetrical, and that does not require further transformations to avoid adverse affects of outliers.
@@
The panels in the middle row present the partial effect of $1-r$ as predictor of RT using FIL (AIC -12900.75), and the bottom panels present the corresponding partial effects using EL (AIC -10896.19). The GAM with the FIL-based predictor clearly provides the superior fit to the observed response latencies. The effect of $1-r$ on the mean is fairly linear for FIL but U-shaped for EL. A strong effect on the variance is present for FIL, but absent for EL. Effects are opposite to those of frequency, as expected, as $1-r$ is constructed to be positively correlated with RT. The absence of a highly significant effect on the variance in RT for EL ($p=0.0356$) is perhaps unsurprising given that EL learns the mapping from form to meaning to perfection (within the constraints of a linear mapping and the type distributions of forms and meanings), and hence little predictivity for human processing variance is to be expected. The question of why $1-r$ has a U-shaped effect on RT for EL will be addressed below.
@@
In summary, FIL generates a predictor ($1-r$) that is better aligned with observed RTs. However, log frequency provides a better fit (AIC -16384.8). Here, it should be kept in mind that word frequency is correlated with many other lexical properties (Baayen,, 2010), including word length and number of senses. Longer, typically less frequent, words require more fixations, and hence are expected to have longer reaction times. The greater number of senses for higher-frequency words are not directly reflected in the embeddings, which typically consist of a single embedding per unique word form, rendering the mapping less precise. As a consequence, there is necessarily imprecision in measures derived from our learning models. Furthermore, $1-r$ is only one of the many learning-based measures that predict lexical decision times, see Heitmeier et al., (2022) for detailed discussion.
@@
Figure 8: Partial effects for mean (left, confidence intervals in blue) and variance (right, confidence intervals in green, the y-axis is on the $\log(\sigma-0.01$ scale), for Gaussian Location-Scale GAMs predicting $1-r$ from log frequency for FIL (upper panels), and for EL (bottom panels). The panel in the upper right zooms in on the partial effect of variance shown to its left. The vertical red lines represent the quartiles of log frequency.
@@ 7.2 Spoken word recognition in Mandarin
It was found that reaction times to the target words are shortest in the ST condition, hence most priming, as expected. Interestingly, mere tone sharing or syllable sharing is not sufficient to induce a reliable priming effect: RTs of both the S and T conditions do not differ significantly from the UR condition. But importantly, there is still a significant difference between the S and T conditions, where syllable sharing induces faster responses than tone sharing. In other words, more priming is found for syllable sharing than tone sharing. It is noteworthy that this pattern of results holds regardless of whether a long (250 ms, experiment 1) or short (50 ms, experiment 2) inter-stimulus interval is adopted in the experiment.
@@ 8 But what about order?
In order to investigate the effect of neglecting order during learning, we also trained a model with the same form and semantic matrix but using the FIL method. The FIL method results in a correlation accuracy of 12.9% for Sarah and 8.0% for Lily. For WHL the correlation accuracies vary across learning rates (Table 3), with accuracy decreasing for lower learning rates. FIL correlation accuracies are somewhat better than WHL accuracies for $\eta=0.001$ and somewhat worse than for $\eta=0.01$. Target correlations obtained with FIL and WHL are in general remarkably similar, correlated the highest for learning rates of $0.001$ and $0.01$ for ``Lily'' and ``Sarah'' respectively. This can also be observed visually in Figure 10: WHL and FIL target correlations are the least similar for $\eta=0.1$ and $\eta=0.0001$. Interestingly, for higher learning rates, low WHL correlations tend to be higher in FIL and vice versa, whereas for lower learning rates an advantage of FIL over WHL is more visible for higher accuracies, while low accuracies in WHL tend to be even lower in FIL.
(a) Sarah (Brown corpus)
(b) Lily (Providence corpus)
@@
Figure 10: Correlation of WHL learned predicted semantics with their targets against correlations of FIL learned predicted semantics with their targets, for different learning rates.
Striking a balance between WHL accuracy and correlation between WHL and FIL accuracies, we now focus on $\eta=0.01$. As can be seen in Figure 10 there are a few outliers where either FIL clearly outperforms WHL or vice versa. For the former case, the most apparent cases for Sarah are no, mummy, cold, later, michael and told. Except no which suffers due to its homophone know, all show a relatively higher frequency at the beginning of the learning trajectory compared to the end (see Figure 10(a)). Moreover, they tend to have overlapping di-phones: for example, told has overlap with old with which they are confused in the WHL model. Lily's data shows a similar pattern. WHL therefore unlearns when two conditions apply: a word occurs very infrequently in the later stages of the learning sequence and it has overlaps with cues in other words. WHL has an advantage over FIL when words have a higher frequency later in the sequence: for Sarah's data this is the case for idea, old, mister, giraffe and alligator (see Figure 10(b)).
(a) FIL over WHL
(b) WHL over FIL
@@
Figure 11: Individual tokens where FIL and WHL correlations to target differ clearly, taken from ``Sarah'', trained with WHL and $\eta=0.01$. (a) shows cases where FIL outperforms WHL, (b) where WHL outperforms FIL. In (a), frequency tends to decrease over time, while in (b) the opposite is the case. Frequencies are normalised by their maximal frequency inside a learning batch of 5000 learning events.
To confirm this qualitative analysis quantitatively, we computed the mean, mode, skewness and kurtosis of the frequency distribution across time for each word. We then used these as predictors in individual (due to collinearity) Generalised Additive Models (GAMs; Wood,, 2011) to predict the difference in target correlation in WHL and FIL for $\eta=0.01$, i.e. predicting whether WHL or FIL perform better depending on the frequency distribution of a word across time (details on method in Supplementary Materials). We found that the higher the mean and mode, and the more negative the skew of the frequency distribution (all indicating that frequencies are higher at later timesteps; compare e.g. ``giraffe'' and ``alligator'' in Figure 11 for examples of high mean and mode and low skewness, with ``mummy'' and ``michael'' in Figure 11 for examples of low mean and mode and high skewness), the better WHL performed than FIL. To illustrate, we show the effect of the mean on performance in the upper row of Figure 12. Additionally, when kurtosis was entered (differentiated by whether skewness was positive or negative), an interesting effect emerged: for negative skewness values (high frequency at later time steps), more positive kurtosis (peakier distribution) yielded an advantage of WHL, while a peakier distribution of positive skewness values lead to an advantage of FIL (see bottom row of Figure 12). All these results confirmed our qualitative analysis above.
@@
Figure 12: Predicting the difference between target correlations in WHL and in FIL from frequency distribution of words across time. Top row: The higher the mean (i.e. higher frequencies at later time steps, see ``giraffe'' and ``alligator'' in the lower row of Figure 11 for examples of words with a high mean), the better WHL performs than FIL. The vertical red lines represent the 0%, 25%, 50%, 75%, and 100% percentiles. Bottom row: For negative skew (higher frequencies at later time steps), the peakier the distribution (higher kurtosis), the larger the advantage of WHL over FIL, and vice versa for positive skew. Kurtosis was transformed to reduce outlier effects, details in Supplementary Materials.

### v2

@@ 3.4.2 High-dimensional modeling
Figure 6: Predicted accuracy@1 as a function of log frequency for high-dimensional representations of form and meaning (red solid line). The light blue dashed line shows the predicted accuracy based on the low-dimensional model, for comparison. The light pink circles represent the target correlations in the high-dimensional model. The small dataset refers to the dataset with 2,638 word forms based on Ernestus and Baayen, (2003), the large dataset to the dataset created from the DLP (Brysbaert and New,, 2009) including 13,669 word forms.
For the small and large datasets, a clear relationship between log frequency and correlation/accuracy is visible.
A model using FIL obtained an average accuracy@1 of 22.3%, which is clearly superior to the 9.9% obtained for the lower-dimensional model. The upper panel of Figure 6 presents the predicted accuracy curves for the high-dimensional FIL model in red (solid line), and the low-dimensional FIL model in blue (dashed line). We see a rise in accuracy for lower frequencies.
3.4.3 Increasing the dataset size
@@
Having established the relationship between EL, WHL and FIL, we can also investigate how EL and FIL fare when the modelled dataset is significantly larger. To this end we used the large dataset introduced above. We found that accuracies in general were clearly lower: For EL, the accuracy@1 was 67.8% and for FIL it was 5.1%. Qualitatively, the differences between EL and FIL are therefore similar (see also Figure 6), but our simple linear mappings clearly perform less well with very large datasets, especially when taking into account frequency.
3.4.4 Frequency-weighted accuracy
@@
A final question is whether the way we have been evaluating accuracy so far is reasonable. In the average accuracy@1 each word's accuracy contributes the same, that is, we have effectively calculated accuracy across all word types in our corpus. However, from a usage-based perspective, comprehending high frequency words is much more important than comprehending low frequency words — this is why, for instance, second language learners are generally taught the most frequent words of a language first. In a usage-based analysis of our model we should therefore be calculating accuracy across word tokens instead. Practically this means that if we go through a corpus of written English, instead of counting how many unique words our model is able to comprehend, we count how many of all of the encountered word tokens are understood correctly. [^7] As pointed out by a reviewer, this accuracy metric puts a lot of emphasis on high frequency closed class words which are often considered stopwords in computational linguistics (examples in English are the, of, and, at, by, of,…; see for instance the English stopword list in NLTK (Bird,, 2006)). However, stopwords have to be learned just as other words, so excluding them a-priori seems unprincipled. We also do not see a plausible cognitive mechanism that would filter out stopwords but not other high frequency open class words. Another issue is that what are considered stopwords in English or Dutch, are often incorporated inside word forms in other languages (e.g. what is expressed with prepositions in English is expressed with case inflection in Finnish). Excluding stopwords for English but including case inflections in Finnish would make modeling incomparable and incompatible across languages.
@@
Following this line of argumentation we also provide frequency-weighted accuracies in Table 1. For instance, the word with the highest frequency in the Dutch Lexicon Project (large dataset) is de (eng. the). Since this word accounts for 7% of all word tokens in the Dutch Lexicon Project (calculated by summing up the frequencies of all word types), it also contributes 7% to our frequency-weighted accuracy measure. There are also 349 words with a frequency of 0 in CELEX (2.6% of all word types) which accordingly do not contribute to the frequency-weighted accuracy at all. We find that with this method, the results flip in comparison with average accuracy@1: Generally, FIL and WHL based on untransformed frequencies perform the best, followed by methods based on log-transformed frequencies, while EL clearly performs the worst. For example, for the high-dimensional simulations on the large dataset, EL has a frequency-weighted accuracy@1 of 43.7%, while FIL achieves 79.8%. This means that EL understands less than half of the word tokens it encounters in a corpus, while FIL comprehends about 8 out of 10 word tokens.
Whether type or token accuracy should be used to evaluate a model depends on the type of analysis conducted by the modeller. For a usage-based perspective as we haven taken here, a token-based measure is more appropriate. For other types of analysis, a type-based measure may be more suited. For instance, in the case of morphologically more complex languages such as Estonian or Russian, the modeller may well be interested in how well the model is able to understand and produce even low-frequency inflected forms of well-known lemmas. A token-based accuracy measure is not helpful in such cases.
@@ 4.1 Visual word recognition in Dutch
In order to assess the possibilities offered by FIL-based modeling (using untransformed frequencies and the high-dimensional form and meaning vectors) for predicting behavioral measures of lexical processing, we return to the large dataset of reaction times to Dutch words described in Section 3.1.1 (13,669 words represented as trigrams with a form dimensionality of 4,678, and semantics represented using 300-dimensional fasttext vectors).
@@
Figure 7: Partial effects for mean (left, confidence intervals in blue) and variance (right, confidence intervals in green, the y-axis is on the $\log(\sigma-0.01)$ scale), for Gaussian Location-Scale GAMs predicting reaction times from log frequency (upper panels), from $1-r$ based on FIL (center panels), and from $1-r$ based on EL (bottom panels). The vertical red lines represent the 0%, 25%, 50%, 75%, and 100% percentiles.
FIL 1-r is a solid predictor for mean and variance in RTs.
Figure 7 presents the partial effects according to three Gaussian Location-Scale GAMs fitted to the response latencies in the Dutch Lexicon Project. Response latencies were inverse transformed (-1000/RT) in order to avoid marked non-normality in the residuals. (Effects for the untransformed RTs are very similar in shape, but confidence intervals are not reliable due to the marked departure from normality of the residuals). The left-hand panels present the partial effects for the mean (in blue), the right panels the partial effects for the variance (in green, on the $\log(\sigma-0.01)$ scale; for further information on how partial effects are calculated in GAMs see Wood, (2017)). The upper panels pertain to a GAM predicting RT from log frequency (AIC -16,384.8; a backoff value of 1 was again added before log-transformation of word frequencies). Mean and variance decrease non-linearly with increasing frequency. The smooth for the mean shows the kind of non-linearity that is typically observed for reaction time data (see, e.g., Baayen,, 2005; Miwa and Baayen,, 2021): the effect of log frequency levels off strongly for high-frequency words and to a lesser extent also for low-frequency words. The partial effect of the variance is less wiggly and decreases as frequency increases. This decrease in variability for increasing frequency has at least two possible sources. First, high-frequency words are known to all speakers, whereas low-frequency words tend to be specialized and known to smaller subsets of speakers. Second, more practice, as in the case of high-frequency words, typically affords reduced variability (see, e.g., Tomaschek et al.,, 2018).
@@
A model-based measure that we expected to correlate with reaction time is the proximity of a word's predicted semantic vector to its corresponding gold standard vector (i.e. the target correlation). The more similar the two vectors are, the better a word's meaning is reconstructed from its form. In other words, the more effective a word form is in approximating its meaning, the more word-like it is and the faster a lexicality decision can be executed. We used the correlation $r$ of $\hat{\mathbf{s}}$ and $\mathbf{s}$ as a measure of semantic proximity. Since for large $r$, lexical decision times are expected to be short, whereas for small $r$ long decision times are more likely, we took $1-r$ as a measure that we expect to enter into a positive correlation with RT. This measure, when based on FIL, has a density that is roughly symmetrical, and that does not require further transformations to avoid adverse affects of outliers.
@@
The panels in the middle row present the partial effect of $1-r$ as predictor of RT using FIL (AIC -12,900.75), and the bottom panels present the corresponding partial effects using EL (AIC -10,896.19). The GAM with the FIL-based predictor clearly provides the superior fit to the observed response latencies. The effect of $1-r$ on the mean is fairly linear for FIL but U-shaped for EL. A strong effect on the variance is present for FIL, but absent for EL. Effects are opposite to those of frequency, as expected, as $1-r$ is constructed to be positively correlated with RT. The absence of a highly significant effect on the variance in RT for EL ($p=0.0356$) is perhaps unsurprising given that EL learns the mapping from form to meaning to perfection (within the constraints of a linear mapping and the type distributions of forms and meanings), and hence little predictivity for human processing variance is to be expected. The question of why $1-r$ has a U-shaped effect on RT for EL will be addressed below.
@@
In summary, FIL generates a predictor ($1-r$) that is better aligned with observed RTs. However, log frequency provides a better fit (AIC -16,384.8). Here, it should be kept in mind that word frequency is correlated with many other lexical properties (Baayen,, 2010), including word length and number of senses. Longer, typically less frequent, words require more fixations, and hence are expected to have longer reaction times. The greater number of senses for higher-frequency words are not directly reflected in the embeddings, which typically consist of a single embedding per unique word form, rendering the mapping less precise. As a consequence, there is necessarily imprecision in measures derived from our learning models. Furthermore, $1-r$ is only one of the many learning-based measures that predict lexical decision times, see Heitmeier et al., 2023a for detailed discussion.
Figure 8: Partial effects for mean (left, confidence intervals in blue) and variance (center and right, confidence intervals in green, the y-axis is on the $\log(\sigma-0.01)$ scale), for Gaussian Location-Scale GAMs predicting $1-r$ from log frequency for FIL (upper panels), and for EL (bottom panels). The panel in the upper right zooms in on the partial effect of variance shown to its left. The vertical red lines represent the quartiles of log frequency.
FIL 1-r shows a similar S-shaped curve as a function of frequency as observed for reaction times, but does not have a similar effect as frequency on the variance in the reaction times.
@@ 4.2 Spoken word recognition in Mandarin
It was found that reaction times to the target words are shortest in the ST condition, hence most priming, as expected. Interestingly, mere tone sharing or syllable sharing is not sufficient to induce a reliable priming effect: the T condition does not differ significantly from the UR condition, whereas the S condition differs from UR only in item analysis, but not in subject analysis. But importantly, there is still a significant difference between the S and T conditions, where syllable sharing induces faster responses than tone sharing. In other words, more priming is found for syllable sharing than tone sharing. It is noteworthy that this pattern of results holds regardless of whether a long (250 ms, experiment 1) or short (50 ms, experiment 2) inter-stimulus interval is adopted in the experiment.
@@ 5 But what about order?
Figure 10: Correlation of WHL learned predicted semantics with their targets against correlations of FIL learned predicted semantics with their targets (blue dots), for different learning rates, and for Sarah (a) and Lily (b). The diagonal red lines denote the x=y line.
WHL target correlations and FIL target correlations are highly correlated, with the tightest relationship visible for $\eta=0.01$ for Sarah and $\eta=0.001$ for Lily (see also Table 3).
In order to investigate the consequences of neglecting order during learning, we also trained a model with the same form and semantic matrix but using the FIL method. The FIL method results in a correlation accuracy of 12.9% for Sarah and 8.1% for Lily. For WHL the correlation accuracies vary across learning rates (Table 3), with accuracy decreasing for lower learning rates. FIL correlation accuracies are somewhat better than WHL accuracies for $\eta=0.001$ and somewhat worse than for $\eta=0.01$. Target correlations obtained with FIL and WHL are in general remarkably similar, correlated the highest for learning rates of $0.001$ and $0.01$ for ``Lily'' and ``Sarah'' respectively. This can also be observed visually in Figure 10: WHL and FIL target correlations are the least similar for $\eta=0.1$ and $\eta=0.0001$. Interestingly, for higher learning rates, low WHL correlations tend to be higher in FIL and vice versa, whereas for lower learning rates an advantage of FIL over WHL is more visible for higher accuracies, while low accuracies in WHL tend to be even lower in FIL.
@@
Figure 11: Individual tokens where FIL and WHL correlations to target differ clearly, taken from ``Sarah'', trained with WHL and $\eta=0.01$. (a) shows cases where FIL outperforms WHL, (b) where WHL outperforms FIL. Frequencies are normalised by their maximal frequency inside a learning batch of 5000 learning events.
In (a), frequency tends to decrease over time. Thus, the item is learned well in the beginning, and the correlation with target goes down over time. In (b)
the opposite is the case.
Figure 11: Individual tokens where FIL and WHL correlations to target differ clearly, taken from ``Sarah'', trained with WHL and $\eta=0.01$. (a) shows cases where FIL outperforms WHL, (b) where WHL outperforms FIL. Frequencies are normalised by their maximal frequency inside a learning batch of 5000 learning events.
In (a), frequency tends to decrease over time. Thus, the item is learned well in the beginning, and the correlation with target goes down over time. In (b)
Striking a balance between WHL accuracy and correlation between WHL and FIL accuracies, we now focus on $\eta=0.01$. As can be seen in Figure 10 there are a few outliers where either FIL clearly outperforms WHL or vice versa. For the former case, the most apparent cases for Sarah are no, mummy, cold, later, michael and told. Except no which suffers due to its homophone know, all show a relatively higher frequency at the beginning of the learning trajectory compared to the end (see upper panel of Figure 11). Moreover, they tend to have overlapping di-phones: for example, told has overlap with old with which they are confused in the WHL model. Lily's data shows a similar pattern. WHL therefore unlearns when two conditions apply: a word occurs very infrequently in the later stages of the learning sequence and it has overlaps with cues in other words. WHL has an advantage over FIL when words have a higher frequency later in the sequence: for Sarah's data this is the case for idea, old, mister, giraffe and alligator (see lower panel of Figure 11).
@@
Figure 12: Predicting the difference between target correlations in WHL and in FIL from frequency distribution of words across time. The y-axes show partial effects. Top row: The higher the mean (i.e. higher frequencies at later time steps, see ``giraffe'' and ``alligator'' in the lower row of Figure 11 for examples of words with a high mean), the better WHL performs than FIL. The vertical red lines represent the 0%, 25%, 50%, 75%, and 100% percentiles. Bottom row: For negative skew (higher frequencies at later time steps), the peakier the distribution (higher kurtosis), the larger the advantage of WHL over FIL, and vice versa for positive skew. Kurtosis was transformed to reduce outlier effects, details in Supplementary Materials.
WHL outperforms FIL for words with higher frequencies at later learning steps.
To confirm this qualitative analysis quantitatively, we computed the mean, mode, skewness and kurtosis of the frequency distribution across time for each word. We then used these as predictors in individual (due to collinearity) Generalised Additive Models (GAMs; Wood,, 2011) to predict the difference in target correlation in WHL and FIL for $\eta=0.01$, i.e. predicting whether WHL or FIL perform better depending on the frequency distribution of a word across time (details on method in Supplementary Materials). We found that the higher the mean and mode, and the more negative the skew of the frequency distribution (all indicating that frequencies are higher at later timesteps; compare e.g. ``giraffe'' and ``alligator'' in Figure 11 for examples of high mean and mode and low skewness, with ``mummy'' and ``michael'' in Figure 11 for examples of low mean and mode and high skewness), the better WHL performed than FIL. To illustrate, we show the effect of the mean on performance in the upper row of Figure 12. Additionally, when kurtosis was entered (differentiated by whether skewness was positive or negative), an interesting effect emerged: for negative skewness values (high frequency at later time steps), more positive kurtosis (peakier distribution) yielded an advantage of WHL, while a peakier distribution of positive skewness values led to an advantage of FIL (see bottom row of Figure 12). All these results confirmed our qualitative analysis above.